# FlashKAT: Understanding and Addressing Performance Bottlenecks in the Kolmogorov-Arnold Transformer

### v1

@@ Abstract
The Kolmogorov-Arnold Network (KAN) has been gaining popularity as an alternative to the multi-layer perceptron (MLP) with its increased expressiveness and interpretability. However, the KAN can be orders of magnitude slower due to its increased computational cost and training instability, limiting its applicability to larger-scale tasks. Recently, the Kolmogorov-Arnold Transformer (KAT) has been proposed, which can achieve FLOPs similar to the traditional Transformer with MLPs by leveraging Group-Rational KAN (GR-KAN). Unfortunately, despite the comparable FLOPs, our characterizations reveal that the KAT is still 123x slower in training speeds, indicating that there are other performance bottlenecks beyond FLOPs. In this paper, we conduct a series of experiments to understand the root cause of the slowdown in KAT. We uncover that the slowdown can be isolated to memory stalls and, more specifically, in the backward pass of GR-KAN caused by inefficient gradient accumulation. To address this memory bottleneck, we propose FlashKAT, which builds on our restructured kernel that minimizes gradient accumulation with atomic adds and accesses to slow memory. Evaluations demonstrate that FlashKAT can achieve a training speedup of 86.5x compared with the state-of-the-art KAT, while reducing rounding errors in the coefficient gradients. Our code is available at https://github.com/OSU-STARLAB/FlashKAT.
@@ 1 Introduction
The Kolmogorov-Arnold Network (KAN) has emerged as a popular alternative to the multi-layer perceptron (MLP) liu2024kan . Rather than learning a fixed weight on each edge, the KAN instead learns a flexible non-linearity. In doing so, the KAN is capable of increased expressivity and interpretability compared to the MLP. Initial works have primarily demonstrated the impressive capabilities of the KAN in scientific applications or equation modeling tasks li2025u ; coffman2025matrixkan ; liu2024kan ; koenig2024kan ; wang2024kolmogorov . However, even with the superior performance of the KAN for these tasks, it has fallen short on its applicability to larger-scale tasks in NLP and Computer Vision due to substantially increased parameters and computational cost, lack of GPU hardware optimization (e.g., for basis functions), and training instability yu2024kan ; le2024exploring ; VisionKAN2024 . For example, due to these non-linear edges, an edge may require up to 204 FLOPs compared to an MLP, which would require only 2 FLOPs per edge (1 for multiply and 1 for add) yang2024kolmogorov . Therefore, significant research is needed to reduce the orders of magnitude performance gap between KAN and MLP.
@@
Recently, the Kolmogorov-Arnold Transformer (KAT) has managed to overcome some of the obstacles preventing the KAN from being deployed to large-scale tasks yang2024kolmogorov . It achieves this by replacing each MLP in the Transformer with a Group-Rational KAN (GR-KAN), which augments the input to each linear layer with the safe Pade Activation Unit (PAU)molina2019pad in a grouped fashion (more details in Section 2). In doing so, each GR-KAN would learn a weighted summation of learned nonlinear functions. Unlike the traditional KAN, the GR-KAN is more computationally efficient, reducing the number of FLOPs to only slightly more than that of MLP, and is constructed to better map to GPUs. By modifying the Transformer with GR-KAN, the resulting KAT is capable of achieving impressive results on computer vision tasks.
@@
However, even with a comparable number of FLOPs (and similar expected speed) offered by the KAT, surprisingly, the KAT turns out to be significantly slower compared to its MLP counterpart in our characterization. For instance, our experiments reveal that a KAT (using GR-KAN) is 123x slower during training than an identically sized Vision Transformer (ViT, using MLP) dosovitskiy2020image on Imagenet-1K russakovsky2015imagenet . This prompts us to conduct a series of experiments to fully understand the performance bottleneck of KAT. When isolating the source of the problem, we have identified that the slowdown is caused by a memory bottleneck rather than excessive computation. More specifically, the source of the bottleneck stems from an inefficient method of accumulating gradients for the coefficients of the PAU, which uses atomic adds.
@@
To address the memory bottleneck, we propose FlashKAT, a substantially faster KAT that builds on our optimized method of GR-KAN that not only avoids accumulating gradients with atomic adds but also avoids excessive accesses to high-bandwidth memory (HBM). The method restructures data accesses and allows us to leverage shared memory to perform a sum reduction for gradient accumulations. By implementing our method as a GPU kernel, the resulting FlashKAT can achieve a training speedup of 86.5x compared with the state-of-the-art KAT. Furthermore, we demonstrate that our new kernel can reduce rounding errors in the gradient computations of the PAU coefficients by nearly an order of magnitude, thus improving training stability.
@@
    Conduct a detailed investigation to understand the computational inefficiencies of the KAT, revealing that the primary performance bottleneck arises not from FLOPs, but from excessive memory traffic caused by inefficient gradient accumulation.
@@
    Propose FlashKAT, which addresses the identified memory bottleneck by leveraging a substantially more efficient gradient accumulation method to speed up training.
@@
    Perform an extensive evaluation of FlashKAT, demonstrating significantly reduced training times and rounding errors from the proposed method.
@@ 2 Background and Related Works
2.1 Kolmogorov-Arnold Network
@@
Recently, the KAN has been popularized as an alternative to the traditional MLP (liu2024kan, ). It is based on the Kolmogorov-Arnold Theorem, which asserts that a multivariate continuous function on a bounded domain $f:[0,1]^{n}\rightarrow\mathbb{R}$ can be rewritten as a finite summation of univariate continuous functions, $\phi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$ as expressed in Equation 1.
$f(x_{1},...,x_{n})=\sum^{2n+1}_{q=1}\phi_{q}(\sum^{n}_{p=1}\phi_{q,p}(x_{p}))$ (1)
@@
Although Equation 1 presents a theoretical form of the Kolmogorov-Arnold Theorem, the authors of (liu2024kan, ) demonstrate its effectiveness in a more generalized form by scaling the width and depth arbitrarily for a defined task. Suppose $n_{l}$ is the number of input nodes for layer $l$, where each input is represented by $x_{l,i}$. Each edge connection to layer $l+1$ has a learned activation $\phi_{l,j,i}(\cdot)$, where $i\in[1,n_{l}]$ and $j\in[1,n_{l+1}]$. Then, each output node is represented by Equation 2.
$x_{l+1,j}=\sum^{n_{l}}_{i=1}\phi_{l,j,i}(x_{l,i})$ (2)
In a vectorized format, we can represent Equation 2 with Equation 3.
@@
$\mathbf{x}_{l+1}=\Phi\circ\mathbf{x}_{l}=\begin{bmatrix}\phi_{l,1,1}(\,\cdot\,)&\cdots&\phi_{l,1,n_{l}}(\,\cdot\,)\\ \vdots&\ddots&\vdots\\ \phi_{l,n_{l+1},1}(\,\cdot\,)&\cdots&\phi_{l,n_{l+1},n_{l}}(\,\cdot\,)\end{bmatrix}\mathbf{x}_{l}$ (3)
Thus, if we have a $L$ layer KAN, it can be represented by Equation 4.
$\text{KAN}(\mathbf{x})=(\Phi_{L}\circ\Phi_{L-1}\circ...\circ\Phi_{1})\mathbf{x}$ (4)
@@
The most widely adopted method for representing $\phi_{l,j,i}(\cdot)$ has been with B-splines liu2024kan , and although the B-spline-based KANs demonstrate promising results, they still have a host of shortcomings yang2024kolmogorov ; yu2024kan . These shortcomings consist of (1) an increased parameter and computational cost compared to the MLP alternative (e.g. increased FLOPs as summarized in Table 1), (2) B-splines not being GPU optimizable (e.g., requiring a recursive algorithm), and (3) training instability in larger networks yang2024kolmogorov . Such shortcomings limit the compatibility of the KAN with the Transformer, restricting its application to larger-scale tasks like computer vision vaswani2017attention .
2.2 Kolmogorov-Arnold Transformer
@@
Unlike its predecessors, littered with shortcomings, the KAT succeeds in merging the KAN into a Transformer. It does so by developing a variation of the KAN called GR-KAN, which (1) shares coefficients across edge groups, reducing FLOPs, (2) leverages safe PAU activation functions, eliminating recursion, and (3) applies a variance-preserving initialization, increasing training stability yang2024kolmogorov . Using the GR-KAN, the authors of yang2024kolmogorov construct the KAT by replacing the MLP with the GR-KAN, where each weight in the MLP is augmented with a non-linear, learned activation function (i.e., a non-linear edge in the KAN). Below are more details.
@@
First, rather than learning $d_{in}\cdot d_{out}$ activations, $\phi_{l,j,i}(\cdot)$, like a KAN, the GR-KAN groups inputs and only learns $\lfloor d_{in}/d_{g}\rfloor$ activations, $F(\cdot)$, where $d_{g}$ is the number of activations that share the same set of learned coefficients in a group. Each output $x_{l+1,j}$ is then a weighted summation of $F_{\lfloor i/d_{g}\rfloor}(x_{l,i})$ using learnable weights, $\mathbf{W}\in\mathbb{R}^{d_{out}\times d_{in}}$. The GR-KAN can be expressed in Equation 5.
[…]
@@
Second, each activation in GR-KAN is a PAU molina2019pad , which we will refer to as a group-wise rational, composed of learnable coefficients $a_{i}$ and $b_{j}$ as shown in Equation 6.
$F(x)=\frac{P(x)}{Q(x)}=\frac{a_{0}+a_{1}x+a_{2}x^{2}+...+a_{m}x^{m}}{1+|b_{1}x+b_{2}x^{2}+...+b_{n}x^{n}|}$ (6)
@@
Third, a variance-preserving initialization is applied for stable learning yang2024kolmogorov . The procedure for applying variance preserving weights requires (1) initializing the coefficients of $F(\cdot)$ to mimic a known activation function (ie. Swish, GeLU, Identity) and (2) initializing the weights in $\mathbf{W}$ according to $\mathcal{N}(0,\frac{\alpha}{d_{in}})$ whereby $\alpha=\frac{\mathbb{E}[\mathbf{F(x)}^{2}]}{Var[\mathbf{x}]}$ with the assumption that $\mathbf{x}\sim\mathcal{N}(0,1)$.
@@
Although the KAT is effective at vision tasks and is more computationally efficient than a ViT using the original KAN for training (a result of overcoming shortcomings (1), (2), and (3)), they are still far short in terms of training speed compared to an identical ViT using an MLP. We demonstrate such results in the next Section.
@@ 3 Beyond FLOPs: Understanding Performance Bottlenecks of KAT
Although the KAT presents a promising alternative to the ViT, it still lags behind the ViT in training time. We illustrate this by comparing the combined forward and backward times for KAT-T (5.7M parameters), KAT-S (22.1M parameters), and KAT-B (86.6M parameters) against identically sized ViT-T, ViT-S, and ViT-B models, using the Imagenet-1k dataset for training russakovsky2015imagenet . For the experiment, each KAT group-wise rational contains 8 groups comprising of 6 numerator and 4 denominator coefficients, and all measurements are taken using an H200 GPU. The results, depicted in Figure 1, show the average time of 100 forward and backward passes, following 5 warmup passes. The data reveals that, for these forward/backward passes, KAT-T is 102x slower than ViT-T, KAT-S is 123x slower than ViT-S, and KAT-B is 116x slower than ViT-B.
@@
It is natural to assume that such a dropoff in training speed between the ViT and KAT results from the difference in FLOPs, as the KAT has more FLOPs than the ViT. However, by comparing the estimated FLOPs between the GR-KAN and MLP (Table 1), the only component in the ViT and KAT where they differ, we can dismiss such assumptions.
Table 1: Comparison of parameter counts and FLOPs for different layers. “FuncFLOPs” denotes the FLOPs of the non-linear activation. In KAN, $K$ is the spline order and $G$ the number of intervals; in GR-KAN, $m,n$ are the polynomial degrees and $g$ the number of groups.
Name Number of parameters FLOPs
MLP (ViT) $d_{\mathrm{in}}\times d_{\mathrm{out}}+d_{\mathrm{out}}$ $\mathrm{FuncFLOPs}\times d_{\mathrm{out}}+2\times d_{\mathrm{in}}\times d_{\mathrm{out}}$
KAN $d_{\mathrm{in}}\times d_{\mathrm{out}}\,(G+K+3)+d_{\mathrm{out}}$
$\mathrm{FuncFLOPs}\times d_{\mathrm{in}}+d_{\mathrm{in}}\times d_{\mathrm{out}}$
×$\times\bigl{[}9K\times(G+1.5K)+2G-2.5K+3\,\bigr{]}$
GR-KAN (KAT) $d_{\mathrm{in}}\times d_{\mathrm{out}}+d_{\mathrm{out}}+(m+n\times g)$ $(2m+2n+3)\times d_{\mathrm{in}}+2\times d_{\mathrm{in}}\times d_{\mathrm{out}}$
As can be seen, between the FLOPs expressions of the ViT and KAT, the main difference resides in the MLP activation requiring $\text{FuncFLOPs}\times d_{out}$ FLOPs and the GR-KAN requiring $(2m+2n+3)\times d_{in}$ FLOPs, both of which are negligible compared with the $2\times d_{in}\times d_{out}$ term. This suggests that the slightly increased FLOPs in the KAT may not be the source of the slowdown. To further verify this, we conduct an experiment that artificially increases the FLOPs inside the group-wise rational in the KAT by placing an additional loop around the floating point computations. This would examine whether FLOPs bottleneck the performance. We measure performance metrics using Nvidia Nsight Compute [^1] The documentation for Nvidia Nsight Compute: https://docs.nvidia.com/nsight-compute/ on an RTX 4060 Ti for the forward and backward pass. The dimension configuration is: input tensor, $\mathbf{X}\in\mathbb{R}^{1024\times 197\times 768}$, numerator coefficients, $\mathbf{A}\in\mathbb{R}^{8\times 6}$, denominator coefficients, $\mathbf{B}\in\mathbb{R}^{8\times 4}$, and the output gradient tensor, $\mathbf{dO}\in\mathbb{R}^{1024\times 197\times 768}$. We report these results in Table 2.
Table 2: Performance breakdown for forward and backward passes of the group-wise rational for the KAT. “Loops” is the total number of instructions loops; “Cycles” is the cycle count; “Time” is wall-clock time; “FLOPs” is the floating point operation count; “SM Thp.”, “L1 Thp.”, “L2 Thp.”, and “HBM Thp.” report the percentage of the device’s peak arithmetic and peak memory bandwidth.
@@
From Table 2, we can immediately see that, even if the FLOPs scale to 8x of the original amount, the overall required "Cycles" or "Time" to complete the computation does not scale in either the forward or backward pass. Such results clearly demonstrate that the forward and backward pass kernels in KAT are not compute-bound, and we must look beyond FLOPs for performance bottlenecks.
@@
Results in Table 2 also demonstrate that a far greater amount of compute time is dedicated to the backward pass than the forward pass. For instance, the forward pass takes 4.96 ms, whereas the backward pass requires 1.03 s, which is 207.7x as long. Therefore, in order to improve the training time of the KAT, performance issues for the backward pass computation must be identified and addressed. The backward pass can be broken down into three components (1) loading the values to compute the gradients, (2) computing the gradients, and (3) storing the gradients. Since we know the backward pass is not compute-bound from Insight 2, this means that (1) and (2) are to blame for the slowdown, both of which are memory issues.
@@
005005005005001,00010001{,}0001 ,  [^0] Mem Stall - Long ScoreboardMem Stall - LG ThrottleMem Stall - MIO ThrottleMem Stall - MembarMem Stall - BarrierMem Stall - DrainMem Stall - Short ScoreboardMem Stall - Tex ThrottleComputing - Selected981.51981.51981.51981.51680.17680.17680.17680.17454.78454.78454.78454.78387.62387.62387.62387.62220.57220.57220.57220.5764.1164.1164.1164.1120.9520.9520.9520.955.715.715.715.712.382.382.382.38Cycles
@@
Figure 2: The warp-state statistics for the KAT group-wise rational backward pass. "Computing - Selected" is where the warp does useful computation; all the others are memory-stall states.

### v2

@@ Abstract
The Kolmogorov-Arnold Network (KAN) has been gaining popularity as an alternative to the multi-layer perceptron (MLP) with its increased expressiveness and interpretability. Even so, the KAN suffers from being orders of magnitude slower due to its increased computational cost and training instability, limiting its applicability to larger-scale tasks. Recently, the Kolmogorov-Arnold Transformer (KAT) has been proposed, which can achieve FLOPs similar to the traditional Transformer with MLPs by leveraging Group-Rational KAN (GR-KAN). Unfortunately, despite the comparable FLOPs, our testing reveals that the KAT is still 123x slower in training speeds, indicating that there are other performance bottlenecks beyond FLOPs. In this paper, we conduct a series of experiments to understand the root cause of the slowdown in KAT. We uncover that the slowdown can be isolated to memory stalls, linked more specifically to inefficient gradient accumulations in the backward pass of GR-KAN. To address this memory bottleneck, we propose FlashKAT, which minimizes accesses to slow memory and the usage of atomic adds through a restructured kernel. Evaluations demonstrate that FlashKAT can achieve a training speedup of 86.5x compared with the state-of-the-art KAT, while reducing rounding errors in the computation of the gradients.
Codehttps://github.com/OSU-STARLAB/FlashKAT
@@ 1 Introduction
The Kolmogorov-Arnold Network (KAN) has emerged as a popular alternative to the multi-layer perceptron (MLP) (liu2024kan). Rather than learning a fixed weight on each edge, the KAN instead learns a flexible nonlinearity. In doing so, the KAN is capable of increased expressivity and interpretability compared to the MLP. Initial works have primarily demonstrated the impressive capabilities of the KAN in scientific applications and equation modeling tasks (li2025u; coffman2025matrixkan; liu2024kan; koenig2024kan; wang2024kolmogorov). However, even with the superior performance of the KAN for these tasks, it has fallen short on larger-scale tasks in NLP and computer vision due to substantially increased parameters and computational cost, lack of GPU hardware optimization (e.g., for basis functions), and training instability (yu2024kan; le2024exploring; VisionKAN2024). For example, while computing an edge for an MLP only requires 2 FLOPS (1 for multiply and 1 for add), computing an edge for KAN may require up to 204 FLOPs, due to the nonlinearity (yang2024kolmogorov). Such a shortcoming necessitates significant research toward reducing the orders of magnitude performance gap between KAN and MLP.
@@
Recently, the Kolmogorov-Arnold Transformer (KAT) has managed to overcome some of the obstacles preventing the KAN from being deployed to large-scale tasks (yang2024kolmogorov). It achieves this by replacing each MLP in the Transformer with a Group-Rational KAN (GR-KAN), which augments the input to each linear layer with the safe Pade Activation Unit (PAU)(molina2019pad) in a grouped fashion (more details in Section 2). In doing so, each GR-KAN would learn a weighted summation of learned nonlinear functions. Unlike the traditional KAN, the GR-KAN is more computationally efficient, reducing the number of FLOPs to only slightly more than that of an MLP, and is constructed to better map to GPUs. By modifying the Transformer with GR-KAN, the resulting KAT is capable of achieving impressive results on computer vision tasks.
@@
However, even with the comparable number of FLOPs (and similar expected speed) offered by the KAT, in our characterization, the KAT is significantly slower compared to its MLP counterpart. For instance, our experiments reveal that a KAT (using GR-KAN) is 123x slower during training than an identically sized Vision Transformer (ViT, using MLP) (dosovitskiy2020image) on ImageNet-1K (russakovsky2015imagenet). This realization motivated an in-depth investigation to comprehensively understand the root causes of KAT’s performance bottleneck. Unlike previous approaches focused primarily on FLOPs, our methodology re-examines the performance issue from a memory-centric perspective, a novel angle not previously explored (yang2024kolmogorov), which uncovers that the slowdown is primarily due to a memory bottleneck rather than excessive computation. Specifically, we identify that the inefficient gradient accumulation method for the PAU coefficients, involving atomic adds, is the key contributor to this bottleneck (molina2019pad). This insight not only challenges assumptions about computation as the main constraint in the KAT but also provides a new direction for optimizing memory access patterns in similar architectures.
@@
To address the memory bottleneck, we propose FlashKAT, a substantially faster KAT that solves the GR-KAN (and by extension the PAU) shortcomings by not only avoiding accumulating gradients with atomic adds but also the unnecessary accesses to high-bandwidth memory (HBM). The method restructures data accesses and allows us to leverage shared memory to perform a sum reduction for gradient accumulations. By implementing our method as a GPU kernel, the resulting FlashKAT can achieve a training speedup of 86.5x compared with the state-of-the-art KAT. Furthermore, we demonstrate that our new kernel can reduce rounding errors in the gradient computations of the PAU coefficients by nearly an order of magnitude, thus improving training stability.
As our method is a direct optimization of the PAU, a type of learnable rational activation function, the techniques and analysis also apply to alternative learnable rational activation functions, a complementary area of research to the KAT (molina2019pad; delfosse2021recurrent; delfosse2020rationals). Despite the clear need for efficient implementations, researchers in both the KAT and rational activation communities had not identified a way to accelerate these activations. FlashKAT provides this essential optimization, representing a nontrivial breakthrough that not only advances state-of-the-art but also inspires future research in the field.
@@
    Conducting a detailed investigation to understand the computational inefficiencies of the KAT, revealing that the primary performance bottleneck arises not from FLOPs, but from excessive memory traffic caused by inefficient gradient accumulation.
@@
    Proposing FlashKAT, which addresses the identified memory bottleneck by leveraging a substantially more efficient gradient accumulation method to speed up training.
@@
    Performing an extensive evaluation of FlashKAT, demonstrating significantly reduced training times and rounding errors from the proposed method.
@@ 2 Background and Related Works
Kolmogorov-Arnold Network
@@
Recently, the KAN has been popularized as an alternative to the traditional MLP (liu2024kan). It is based on the Kolmogorov-Arnold Theorem (kolmogorov1961representation), which asserts that a multivariate continuous function on a bounded domain $f:[0,1]^{n}\rightarrow\mathbb{R}$ can be rewritten as a finite summation of univariate continuous functions, $\phi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$ as expressed in Equation 1:
$f(x_{1},...,x_{n})=\sum^{2n+1}_{q=1}\phi_{q}(\sum^{n}_{p=1}\phi_{q,p}(x_{p})).$ (1)
@@
Although Equation 1 presents a theoretical form of the Kolmogorov-Arnold Theorem, the authors of (liu2024kan) demonstrated its effectiveness in a more generalized form by scaling the width and depth arbitrarily for a defined task. Suppose $n_{l}$ is the number of input nodes for layer $l$, where each input is represented by $x_{l,i}$. Each edge connection to layer $l+1$ has a learned activation $\phi_{l,j,i}(\cdot)$, where $i\in[1,n_{l}]$ and $j\in[1,n_{l+1}]$. Then, we can represent each output node as
$x_{l+1,j}=\sum^{n_{l}}_{i=1}\phi_{l,j,i}(x_{l,i}).$ (2)
In a vectorized format, we can represent Equation 2 with
@@
$\mathbf{x}_{l+1}=\Phi\circ\mathbf{x}_{l}=\begin{bmatrix}\phi_{l,1,1}(\,\cdot\,)&\cdots&\phi_{l,1,n_{l}}(\,\cdot\,)\\ \vdots&\ddots&\vdots\\ \phi_{l,n_{l+1},1}(\,\cdot\,)&\cdots&\phi_{l,n_{l+1},n_{l}}(\,\cdot\,)\end{bmatrix}\mathbf{x}_{l}.$ (3)
Thus, if we have a $L$ layer KAN, the KAN can be represented as
$\text{KAN}(\mathbf{x})=(\Phi_{L}\circ\Phi_{L-1}\circ...\circ\Phi_{1})\mathbf{x}.$ (4)
@@
The most widely adopted method for representing $\phi_{l,j,i}(\cdot)$ has been with B-splines (liu2024kan), and although the B-spline-based KANs demonstrate promising results, they still have a host of shortcomings (yang2024kolmogorov; yu2024kan). These shortcomings consist of (1) an increased parameter and computational cost compared to the MLP alternative (e.g. increased FLOPs as summarized in Table 1), (2) B-splines not being GPU optimizable (e.g., requiring a recursive algorithm), and (3) training instability in larger networks (yang2024kolmogorov). Such shortcomings limit the KAN’s compatibility with the Transformer, restricting its application to larger-scale tasks like computer vision (vaswani2017attention).
Kolmogorov-Arnold Transformer
@@
Unlike its predecessors, littered with shortcomings, the KAT succeeds in merging the KAN into a Transformer. It does so by developing a variation of the KAN called GR-KAN, which (1) shares coefficients across edge groups, reducing FLOPs, (2) leverages safe PAU activation functions, eliminating recursion, and (3) applies a variance-preserving initialization, increasing training stability (yang2024kolmogorov). Using the GR-KAN, the authors of (yang2024kolmogorov) construct the KAT by replacing the MLP with the GR-KAN, where each weight in the MLP is augmented with a nonlinear, learned activation function (i.e., a nonlinear edge in the KAN).
Table 1: Comparison of parameter counts and FLOPs for different layers. “FuncFLOPs” denotes the activation FLOPs. In KAN, $K$ is the spline order and $G$ the number of intervals; in GR‐KAN, $m,n$ are the polynomial degrees and $g$ the number of groups.
Name Number of parameters FLOPs
MLP (ViT) $d_{\mathrm{in}}\times d_{\mathrm{out}}$ $\mathrm{FuncFLOPs}\times d_{\mathrm{out}}+2\times d_{\mathrm{in}}\times d_{\mathrm{out}}$
KAN $d_{\mathrm{in}}\times d_{\mathrm{out}}\times(G+K+3)$
$\mathrm{FuncFLOPs}\times d_{\mathrm{in}}+d_{\mathrm{in}}\times d_{\mathrm{out}}\times\bigl[9K\times(G+1.5K)+2G-2.5K+3\,\bigr]$
GR‐KAN (KAT) $d_{\mathrm{in}}\times d_{\mathrm{out}}+(m+n\times g+1)$ $(2m+2n+3)\times d_{\mathrm{in}}+2\times d_{\mathrm{in}}\times d_{\mathrm{out}}$
@@
First, rather than learning $d_{in}\cdot d_{out}$ activations, $\phi_{l,j,i}(\cdot)$, like a KAN, the GR-KAN groups inputs and only learns $\lfloor d_{in}/d_{g}\rfloor$ activations, $F(\cdot)$, where $d_{g}$ is the number of activations that share the same set of learned coefficients in a group. Each output $x_{l+1,j}$ is then a weighted summation of $F_{\lfloor i/d_{g}\rfloor}(x_{l,i})$ using learnable weights, $\mathbf{W}\in\mathbb{R}^{d_{out}\times d_{in}}$. The GR-KAN is expressed in Equation 5:
$\begin{split}&\text{GR-KAN}(\mathbf{x})=\,\mathbf{W}\mathbf{F}(\mathbf{x})\\ &=\begin{bmatrix}w_{11}&\cdots&w_{1,d_{in}}\\ \vdots&\ddots&\vdots\\ w_{d_{out},1}&\cdots&w_{d_{out},d_{in}}\end{bmatrix}\begin{bmatrix}F_{\lfloor 1/d_{g}\rfloor}(x_{1})\\ \vdots\\ F_{\lfloor d_{in}/d_{g}\rfloor}(x_{d_{in}})\end{bmatrix}\end{split}$. (5)
@@
Second, each activation in GR-KAN is a PAU (molina2019pad), which we will refer to as a group-wise rational function, composed of learnable coefficients $a_{i}$ and $b_{j}$ as follows:
$F(x)=\frac{P(x)}{Q(x)}=\frac{a_{0}+a_{1}x+a_{2}x^{2}+...+a_{m}x^{m}}{1+|b_{1}x+b_{2}x^{2}+...+b_{n}x^{n}|}.$ (6)
@@
Third, a variance-preserving initialization is applied for stable learning (yang2024kolmogorov). The procedure for applying variance preserving weights requires (1) initializing the coefficients of $F(\cdot)$ to mimic a known activation function (ie. Swish, GeLU, Identity) and (2) initializing the weights in $\mathbf{W}$ according to $\mathcal{N}(0,\frac{\alpha}{d_{in}})$ whereby $\alpha=\frac{\mathbb{E}[\mathbf{F(x)}^{2}]}{Var[\mathbf{x}]}$ assuming $\mathbf{x}\sim\mathcal{N}(0,1)$.
@@
Although the KAT is effective at vision tasks and is more computationally efficient than a ViT using the original KAN for training (a result of overcoming shortcomings (1), (2), and (3)), they still far short in terms of training speed compared to an identical ViT using an MLP, a shortcoming which has slowed further research. We demonstrate such results in the next section.
@@ 3 Beyond FLOPs: Understanding Performance Bottlenecks of KAT
Although the KAT presents a promising alternative to the ViT, it still lags behind the ViT in training time. We illustrate this by comparing the combined forward and backward times for KAT-T (5.7M parameters), KAT-S (22.1M parameters), and KAT-B (86.6M parameters) against identically sized ViT-T, ViT-S, and ViT-B models, using the ImageNet-1k dataset for training (russakovsky2015imagenet). For the experiment, each KAT group-wise rational function contains 8 groups comprising of 6 numerator and 4 denominator coefficients, and all measurements are taken using an H200 GPU. The results, depicted in Figure 1, show the average time of 100 forward and backward passes, following 5 warmup passes. The data reveals that, for these forward/backward passes, KAT-T is 102x slower than ViT-T, KAT-S is 123x slower than ViT-S, and KAT-B is 116x slower than ViT-B.
@@
It is natural to assume that such a dropoff in training speed between the ViT and KAT results from the difference in FLOPs, as the KAT has more FLOPs than the ViT. However, by comparing the estimated FLOPs between the GR-KAN and MLP (Table 1), the only component in the ViT and KAT where they differ, we can dismiss such an assumption.
As can be seen, between the FLOPs expressions of the ViT and KAT, the main difference resides in the MLP activation requiring $\text{FuncFLOPs}\times d_{out}$ FLOPs and the GR-KAN requiring $(2m+2n+3)\times d_{in}$ FLOPs, both of which are negligible compared with the $2\times d_{in}\times d_{out}$ term. This suggests that the slightly increased FLOPs in the KAT may not be the source of the slowdown. To further verify this, we conduct an experiment that artificially increases the FLOPs inside the group-wise rational function in the KAT by placing an additional loop around the floating point computations. This would examine whether FLOPs bottleneck the performance. We measure performance metrics using Nvidia Nsight Compute [^1] The documentation for Nvidia Nsight Compute can be found at https://docs.nvidia.com/nsight-compute/ on an RTX 4060 Ti for the forward and backward pass. The dimension configuration is: input tensor, $\mathbf{X}\in\mathbb{R}^{1024\times 197\times 768}$, numerator coefficients, $\mathbf{A}\in\mathbb{R}^{8\times 6}$, denominator coefficients, $\mathbf{B}\in\mathbb{R}^{8\times 4}$, and the output gradient tensor, $\mathbf{dO}\in\mathbb{R}^{1024\times 197\times 768}$. We report these results in Table 2.
Table 2: Performance breakdown for forward and backward passes of the group-wise rational function. “Loops” is the total instruction loops; “Cycles” is the cycle count; “Time” is wall-clock time; “FLOPs” is the floating point operation count; “SM Thp.”, “L1 Thp.”, “L2 Thp.”, and “HBM Thp.” report the percentage of the device’s peak arithmetic and memory throughput.
@@
From Table 2, we can immediately see that, even if the FLOPs scale to 8x of the original amount, the overall required ”Cycles” or ”Time” to complete the computation does not scale in either the forward or backward pass. Such results clearly demonstrate that the forward and backward pass kernels in KAT are not compute-bound, and we must look beyond FLOPs for performance bottlenecks.
@@
Results in Table 2 also demonstrate that the group-wise rational function dedicates a far greater amount of compute time to the backward pass than the forward pass. For instance, the forward pass takes 4.96 ms, whereas the backward pass requires 1.03 s, which is 207.7x as long. Therefore, to improve the training time of the KAT, performance issues for the backward pass must be identified and addressed. The backward pass can be broken down into three components: (1) loading values to compute the gradients, (2) computing the gradients, and (3) storing the gradients. Since we know the backward pass is not compute-bound from Insight 2, this means that (1) and (2) are to blame for the slowdown, both of which are memory issues.
@@
05005001,0001{,} [^0] MS - Long ScoreboardMS - LG ThrottleMS - MIO ThrottleMS - MembarMS - BarrierMS - DrainMS - Short ScoreboardMS - Tex ThrottleComputing - Selected981.51981.51680.17680.17454.78454.78387.62387.62220.57220.5764.1164.1120.9520.955.715.712.382.38Cycles
@@
Figure 2: The warp‐state statistics for the KAT group-wise rational function backward pass. ”Computing - Selected” is where the warp does useful computation; all the others are memory stall (MS) states.